-3 1/4 - (1/2) need help with this


Answer by richwmiller(17219)

This has very little to do with proportions.
convert -1/2 to -2/4
-3 1/4-(-2/4)
convert -3 1/4 to improper fraction -13/4
remember -(-2/4)=+2/4
-13/4+2/4=-11/4=-2 3/4